OpenStudy (anonymous):

How would I Calculate the number of molecules present in each of the following samples. 6.55 g of carbon monoxide. ___________molecules

OpenStudy (anonymous):

Someone please help! :(

OpenStudy (abb0t):

@samjordon

OpenStudy (4n1m0s1ty):

You first need to find the molecular mass of carbon dioxide, You can do this by adding the atomic weights of carbon and oxygen. Once you do that you can convert to moles by dividing the sample mass by the molecular mass.

OpenStudy (anonymous):

never mind... It took me a while but finally got it! thanks for the help

OpenStudy (4n1m0s1ty):

ok great

OpenStudy (anonymous):

i need help with another D:

OpenStudy (anonymous):

can you help me with that??

OpenStudy (4n1m0s1ty):

sure

OpenStudy (anonymous):

k its Calculate the mass of 1.45 ✕ 1025 molecules of SO3. Use a molar mass with at least as many significant figures as the data given.

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

mass of 1.45 x 10^25

OpenStudy (4n1m0s1ty):

ok, so you first need to calculate the molecular mass of SO3. Can you do that?

OpenStudy (anonymous):

yeah the mass of SO3 is 80.06

OpenStudy (4n1m0s1ty):

Ok, so now that you have that we do some dimensional analysis. The problem gives us number of molecules, but we want the mass (grams). In stochiometry problems, the best way to convert between different values is to convert to moles first, once have the number of moles solving for anything else is fairly trivial.

OpenStudy (anonymous):

okay

OpenStudy (4n1m0s1ty):

So yo convert to moles, we know that 1 mole of any substance is equivalent to 6.022 x 10^23 units. In this case its equivalent to 6.022 x 10^23 molecules.

OpenStudy (4n1m0s1ty):

to*

OpenStudy (anonymous):

okay

OpenStudy (4n1m0s1ty):

So since we know that, we just follow the math: \[\frac{1.45 * 10^{25} molecules}{} * \frac{1 mole}{6.022 * 10^{23} molecules } = ?\] The molcule units cancel out and we are left with the number of moles. I left the arithmetic as an exercise for you to solve, but do you understand up to this point?

OpenStudy (anonymous):

yeah i do i see where i made my mistake... haha i put the 6.02 x 10^23 ontop :/

OpenStudy (4n1m0s1ty):

good, so you know how to solve it from here? you just take moles and convert to mass using the molecular mass from before.

OpenStudy (anonymous):

wait what? lol im so dumb i got confused again :(

OpenStudy (4n1m0s1ty):

yeah sorry , I'm switching between 2 questions. Ok, so now that you have the number of moles of SO3, you setup another fraction. \[\frac{?\,moles\,SO3}{}*\frac{80.06\,g\,SO3}{1\,mole\,SO3} = ??\,g\,SO3\] where ? is what you just solved for, and the ?? is the result. I just setup the 2nd fraction to eliminate the units I don't want on the bottom and the desired unit on top.

OpenStudy (anonymous):

haha your good

OpenStudy (4n1m0s1ty):

Thanks

OpenStudy (anonymous):

i still got it wrong :(

OpenStudy (anonymous):

so i solved the first equation on top and got 2.41. I then used that number to plug in the ? on the second equation and did 2.41 x 80.06. The final answer was than 192, but it wasnt correct

OpenStudy (4n1m0s1ty):

hold on let me check your work

OpenStudy (4n1m0s1ty):

I got 24.08 moles for the first step. I would check the math again.

OpenStudy (anonymous):

oh woah okay

OpenStudy (4n1m0s1ty):

Did you get it?

OpenStudy (anonymous):

yeah! Thank you so much! you helped :)

OpenStudy (4n1m0s1ty):

no problem